# ORCID: $0$-0002-6512-4229 (Om Prakash)

### v1

@@ 5 Constructions of quantum codes and comparison with the existing codes
Proof
Let $h_{i}^{\prime}(x)h_{i}^{\prime}(x)$ be divisible by $x^{n}-1$ from right for $i=0,1,\dots,s$. Then from Theorem 4.4, we have $\mathcal{C}^{\perp}\subseteq\mathcal{C}$. Also, by Lemma 1, we have $\varphi(\mathcal{C}^{\perp})=\varphi(\mathcal{C})^{\perp}$, and hence $\varphi(\mathcal{C})^{\perp}\subseteq\varphi(\mathcal{C})$. Thus, $\varphi(\mathcal{C})$ is a dual containing linear code with parameters $[(s+1)n,k,d_{H}]$ over $\mathbb{F}_{q}$. Further, by Lemma 2, there exists a quantum code with parameters $[[(s+1)n,2k-(s+1)n,d_{H}]]_{q}$.
Next, with the help of our established results, we construct many new quantum codes possessing better parameters than the existing codes, which are appeared in Edel ; Verma21 . In the following examples, $\mathbb{F}_{q}^{*}=\langle w\rangle$ denotes the cyclic group of non-zero elements of $\mathbb{F}_{q}$ generated by $w\in\mathbb{F}_{q}$. All examples’ computations are carried out using the Magma computation system Bosma .
Example 2
@@
From above, we see that $h_{i}^{\prime}(x)h_{i}^{\prime}(x)$ is divisible by $(x^{30}-1)$ on the right for $i=0,1,2,3$. Hence, by Theorem 5.2, there exists a quantum code with parameters $[[120,108,4]]_{8}$ which has the same length and distance but better code rate than the best-known code $[[120,104,4]]_{8}$ given by Edel .
Example 3
@@
From above we see that $h_{i}^{\prime}(x)h_{i}^{\prime}(x)$ is divisible by $x^{20}-1$ on the right for $i=0,1,2,3$. Hence, by Theorem 5.2, there exists a quantum code with parameters $[[80,68,4]]_{25}$ which has the same length and distance, but better code rate than the best-known code $[[80,64,4]]_{25}$ given by Verma21 .
Let $\mathcal{C}$ be a $(\theta,\Im)$-cyclic code of length $n$ over $\mathbb{F}_{q}$ where $\mathcal{C}=\langle g(x)\rangle$ and $x^{n}-1=h(x)g(x)=g(x)h^{\prime}(x)$ for some monic skew polynomials $g(x),h(x),h^{\prime}(x)\in\mathbb{F}_{q}[x;\theta,\Im]$. Further, let $h^{\prime}(x)h^{\prime}(x)$ be divisible by $x^{n}-1$ from the right. Therefore, by Theorem 4.4, we get the dual containing codes with $\mathbb{F}_{q}$-parameters $[n,k,d]_{q}$ (enlisted in the fourth column of Table 1). Also, by Lemma 2, we construct quantum codes $[[n,k,d]]_{q}$ (in the fifth column), in which some codes satisfy the equality $n-k+2-2d=2$ (Near to MDS), and some are MDS (maximum-distance-separable). Let $\mathcal{C}=\displaystyle\bigoplus_{i=0}^{s}\zeta_{i}\mathcal{C}_{i}$ be a $(\gamma,\Delta)$-cyclic code of length $n$ over $\mathscr{R}_{q,s}$ where $\mathcal{C}_{i}=\langle g_{i}(x)\rangle$ is a $(\theta,\Im)$-cyclic code of length $n$ over $\mathbb{F}_{q}$ and $x^{n}-1=h_{i}(x)g_{i}(x)=g_{i}(x)h^{\prime}_{i}(x)$ for some monic skew polynomials $g_{i}(x),h_{i}(x),h^{\prime}_{i}(x)\in\mathbb{F}_{q}[x;\theta,\Im]$ for $i=0,1,2,\dots,s$. Further, let $h_{i}^{\prime}(x)h_{i}^{\prime}(x)$ is divisible by $x^{n}-1$ from the right for $i=0,1,2,\dots,s$. Therefore, by Theorem 4.4, we get the dual containing codes with $\mathbb{F}_{q}$-parameters $[n,k,d]_{q}$ (enlisted in the fifth column of Table 2). Also, by Theorem 5.2, we construct quantum codes $[[n,k,d]]_{q}$ (in the sixth column), which beat the parameters of best-known codes (in the seventh column) given by the online database Edel ; Verma21 . Also, the first and second columns represent $s$ and $(n,q)$, respectively. Moreover, in third column we present $\theta$-derivations $\Im(a)$ for $a\in\mathbb{F}_{q}$. Note that in fourth column we give generator polynomials $g_{i}$ for $\mathcal{C}_{i}$ ($i=0,1,2,\dots,s$) which is a right factor of $x^{n}-1$ in $\mathbb{F}_{q}[x;\theta,\Im]$. In order to make Table 2 precise, we enlist the coefficients of polynomials in decreasing powers of $x$. For example, we write $w^{7}02w$ to represent the polynomial $w^{7}x^{3}+2x+w$.
Table 1: MDS quantum codes from $(\theta,\Im)$-cyclic codes over $\mathbb{F}_{q}$
$(n,q)$ $\Im(a)$, $a\in\mathbb{F}_{q}$ $g(x)$ $\mathcal{C}$ Obtained Remark
Codes
$(30,8)$ $w(\theta(a)-a)$ $(w^{3}w^{5})$ $[30,29,2]_{8}$ $[[30,28,2]]_{8}$ MDS
$(30,8)$ $w(\theta(a)-a)$ $(ww^{4}w^{6})$ $[30,28,2]_{8}$ $[[30,26,2]]_{8}$ Near to MDS
$(14,49)$ $w^{2}(\theta(a)-a)$ $(w^{39}w^{3}w^{17})$ $[14,12,2]_{49}$ $[[14,10,2]]_{49}$ Near to MDS
$(24,49)$ $w^{2}(\theta(a)-a)$ $(w^{4}1)$ $[24,23,2]_{49}$ $[[24,22,2]]_{49}$ MDS
$(24,49)$ $\theta(a)-a$ $(w^{2}w^{17})$ $[24,23,2]_{49}$ $[[24,22,2]]_{49}$ MDS
$(20,25)$ $w(\theta(a)-a)$ $(w^{14}w^{19}w^{15})$ $[20,18,2]_{25}$ $[[20,16,2]]_{25}$ Near to MDS
$(24,25)$ $\theta(a)-a$ $(w^{3}w^{4}w^{20})$ $[24,22,2]_{25}$ $[[24,20,2]]_{25}$ Near to MDS
$(20,25)$ $w(\theta(a)-a)$ $(ww^{17})$ $[20,19,2]_{25}$ $[[20,18,2]]_{25}$ MDS
@@
$2$ $(48,9)$ $w^{2}(\theta(a)-a)$ $(w^{7}1w^{3},w^{5}w^{2},w^{5}12)$ $[144,139,3]_{9}$ $[[144,134,3]]_{9}$ $[[146,134,3]]_{9}$ Edel
$3$ $(36,9)$ $w^{2}(\theta(a)-a)$ $(w^{2}1w^{5},1w^{3},w^{7}w^{2},2ww^{3})$ $[144,138,4]_{9}$ $[[144,132,4]]_{9}$ $[[146,128,4]]_{9}$ Edel
$3$ $(32,9)$ $w^{2}(\theta(a)-a)$ $(w^{2}ww,w^{6}w^{2}2,2w^{2}ww^{3},w^{3}1)$ $[128,120,4]_{9}$ $[[128,112,4]]_{9}$ $[[129,103,4]]_{9}$ Edel
$2$ $(42,9)$ $w^{2}(\theta(a)-a)$ $(ww^{6},w^{6}ww^{3},w^{7}w^{2})$ $[126,122,3]_{9}$ $[[126,118,3]]_{9}$ $[[130,118,3]]_{9}$ Verma21
$2$ $(60,4)$ $w(\theta(a)-a)$ $(www^{2},1w^{2}w^{2},11w^{2})$ $[180,174,3]_{4}$ $[[180,168,3]]_{4}$ $[[185,167,3]]_{4}$ Verma21
$3$ $(20,25)$ $w(\theta(a)-a)$ $(ww^{17},w^{10}2w^{11},w^{5}3,w^{14}w^{19}w^{15})$ $[80,74,4]_{25}$ $[[80,68,4]]_{25}$ $[[80,64,4]]_{25}$ Verma21
$3$ $(40,25)$ $w(\theta(a)-a)$ $(w^{19}w^{10},w^{10}w^{14},w^{11}w^{17},w^{14}w^{4})$ $[120,116,3]_{25}$ $[[120,112,3]]_{25}$ $[[120,106,3]]_{25}$ Verma21
$3$ $(30,8)$ $w(\theta(a)-a)$ $(w^{2}1,ww^{4}w^{6},w^{4}w^{3}w,1w^{2}w^{4})$ $[120,114,4]_{8}$ $[[120,108,4]]_{8}$ $[[120,104,4]]_{8}$ Edel
$3$ $(32,8)$ $w(\theta(a)-a)$ $(w^{6}w^{3}w^{4},w^{2}w^{5}w^{5},ww^{6},w^{6}w^{5}w^{2})$ $[128,121,4]_{8}$ $[[128,114,4]]_{8}$ $[[128,112,4]]_{8}$ Edel
@@ 6 Conclusion
In this paper, we have constructed many quantum codes over a class of finite commutative non-chain rings $\mathscr{R}_{q,s}$, with better parameters than the codes available in recent literature. Particularly, we have obtained $(\gamma,\Delta)$-cyclic codes using a set of idempotents over $\mathscr{R}_{q,s}$ and established results on their algebraic structure. Towards construction of quantum codes, a necessary and sufficient condition to contain their dual codes has established. Finally, we have obtained many better quantum codes. On the other hand, exploring applications in the quantum computation of these codes is still open as future research work.
@@ Acknowledgements
The first and second authors are thankful to the Department of Science and Technology (DST), Govt. of India, for financial support under CRG/2020/005927, vide Diary No. SERB/F/6780/ 2020-2021 dated 31 December 2020 and Ref No. DST/INSPIRE/03/2016/
001445, respectively.
@@
Data Availability Statement: The authors declare that [the/all other] data supporting the findings of this study are available within the article. Any clarification may be requested from corresponding author provided it is essential.

### v2

@@ 5.2 CSS Code Framework
Proof.
Let $h_{i}^{\prime}(x)h_{i}^{\prime}(x)$ be divisible by $x^{n}-1$ from right for $i=0,1,\dots,s$. Then from Theorem 4.4, we have $\mathcal{C}^{\perp}\subseteq\mathcal{C}$. Also, by Lemma 5.1, we have $\varphi(\mathcal{C}^{\perp})=\varphi(\mathcal{C})^{\perp}$, and hence $\varphi(\mathcal{C})^{\perp}\subseteq\varphi(\mathcal{C})$. Thus, $\varphi(\mathcal{C})$ is a dual containing linear code with parameters $[(s+1)n,k,d_{H}]$ over $\mathbb{F}_{q}$. Further, by Lemma 5.2, there exists a quantum code with parameters $[[(s+1)n,2k-(s+1)n,d_{H}]]_{q}$. ∎
Next, with the help of our established results, we construct many new quantum codes possessing better parameters than the existing codes, which are appeared in [14, 46]. In the following examples, $\mathbb{F}_{q}^{*}=\langle w\rangle$ denotes the cyclic group of non-zero elements of $\mathbb{F}_{q}$ generated by $w\in\mathbb{F}_{q}$. All examples’ computations are carried out using the Magma computation system [10].
Example 5.1.
@@
From above, we see that $h_{i}^{\prime}(x)h_{i}^{\prime}(x)$ is divisible by $(x^{30}-1)$ on the right for $i=0,1,2,3$. Hence, by Theorem 5.2, there exists a quantum code with parameters $[[120,108,4]]_{8}$ which has the same length and distance but better code rate than the best-known code $[[120,104,4]]_{8}$ given by [14].
Example 5.2.
@@
From above we see that $h_{i}^{\prime}(x)h_{i}^{\prime}(x)$ is divisible by $x^{20}-1$ on the right for $i=0,1,2,3$. Hence, by Theorem 5.2, there exists a quantum code with parameters $[[80,68,4]]_{25}$ which has the same length and distance, but better code rate than the best-known code $[[80,64,4]]_{25}$ given by [46].
Let $\mathcal{C}$ be a $(\theta,\Im)$-cyclic code of length $n$ over $\mathbb{F}_{q}$ where $\mathcal{C}=\langle g(x)\rangle$ and $x^{n}-1=h(x)g(x)=g(x)h^{\prime}(x)$ for some monic skew polynomials $g(x),h(x),h^{\prime}(x)\in\mathbb{F}_{q}[x;\theta,\Im]$. Further, let $h^{\prime}(x)h^{\prime}(x)$ be divisible by $x^{n}-1$ from the right. Therefore, by Theorem 4.4, we get the dual containing codes with $\mathbb{F}_{q}$-parameters $[n,k,d]_{q}$ (enlisted in the fourth column of Table 1).
@@
$2$ $(48,9)$ $w^{2}(\theta(a)-a)$ $(w^{7}1w^{3},w^{5}w^{2},w^{5}12)$ $[144,139,3]_{9}$ $[[144,134,3]]_{9}$ $[[146,134,3]]_{9}$ [14]
$3$ $(36,9)$ $w^{2}(\theta(a)-a)$ $(w^{2}1w^{5},1w^{3},w^{7}w^{2},2ww^{3})$ $[144,138,4]_{9}$ $[[144,132,4]]_{9}$ $[[146,128,4]]_{9}$ [14]
$3$ $(32,9)$ $w^{2}(\theta(a)-a)$ $(w^{2}ww,w^{6}w^{2}2,2w^{2}ww^{3},w^{3}1)$ $[128,120,4]_{9}$ $[[128,112,4]]_{9}$ $[[129,103,4]]_{9}$ [14]
$2$ $(42,9)$ $w^{2}(\theta(a)-a)$ $(ww^{6},w^{6}ww^{3},w^{7}w^{2})$ $[126,122,3]_{9}$ $[[126,118,3]]_{9}$ $[[130,118,3]]_{9}$ [46]
$2$ $(60,4)$ $w(\theta(a)-a)$ $(www^{2},1w^{2}w^{2},11w^{2})$ $[180,174,3]_{4}$ $[[180,168,3]]_{4}$ $[[185,167,3]]_{4}$ [46]
$3$ $(20,25)$ $w(\theta(a)-a)$ $(ww^{17},w^{10}2w^{11},w^{5}3,w^{14}w^{19}w^{15})$ $[80,74,4]_{25}$ $[[80,68,4]]_{25}$ $[[80,64,4]]_{25}$ [46]
$3$ $(40,25)$ $w(\theta(a)-a)$ $(w^{19}w^{10},w^{10}w^{14},w^{11}w^{17},w^{14}w^{4})$ $[120,116,3]_{25}$ $[[120,112,3]]_{25}$ $[[120,106,3]]_{25}$ [46]
$3$ $(30,8)$ $w(\theta(a)-a)$ $(w^{2}1,ww^{4}w^{6},w^{4}w^{3}w,1w^{2}w^{4})$ $[120,114,4]_{8}$ $[[120,108,4]]_{8}$ $[[120,104,4]]_{8}$ [14]
$3$ $(32,8)$ $w(\theta(a)-a)$ $(w^{6}w^{3}w^{4},w^{2}w^{5}w^{5},ww^{6},w^{6}w^{5}w^{2})$ $[128,121,4]_{8}$ $[[128,114,4]]_{8}$ $[[128,112,4]]_{8}$ [14]
Also, by Lemma 5.2, we construct quantum codes $[[n,k,d]]_{q}$ (in the fifth column), in which some codes satisfy the equality $n-k+2-2d=2$ (Near to MDS), and some are MDS (maximum-distance-separable). Let $\mathcal{C}=\displaystyle\bigoplus_{i=0}^{s}\zeta_{i}\mathcal{C}_{i}$ be a $(\gamma,\Delta)$-cyclic code of length $n$ over $\mathscr{R}_{q,s}$ where $\mathcal{C}_{i}=\langle g_{i}(x)\rangle$ is a $(\theta,\Im)$-cyclic code of length $n$ over $\mathbb{F}_{q}$ and $x^{n}-1=h_{i}(x)g_{i}(x)=g_{i}(x)h^{\prime}_{i}(x)$ for some monic skew polynomials $g_{i}(x),h_{i}(x),h^{\prime}_{i}(x)\in\mathbb{F}_{q}[x;\theta,\Im]$ for $i=0,1,2,\dots,s$. Further, let $h_{i}^{\prime}(x)h_{i}^{\prime}(x)$ is divisible by $x^{n}-1$ from the right for $i=0,1,2,\dots,s$. Therefore, by Theorem 4.4, we get the dual containing codes with $\mathbb{F}_{q}$-parameters $[n,k,d]_{q}$ (enlisted in the fifth column of Table 1). Also, by Theorem 5.2, we construct quantum codes $[[n,k,d]]_{q}$ (in the sixth column), which beat the parameters of best-known codes (in the seventh column) given by the online database [14, 46]. Also, the first and second columns represent $s$ and $(n,q)$, respectively. Moreover, in third column we present $\theta$-derivations $\Im(a)$ for $a\in\mathbb{F}_{q}$. Note that in fourth column we give generator polynomials $g_{i}$ for $\mathcal{C}_{i}$ ($i=0,1,2,\dots,s$) which is a right factor of $x^{n}-1$ in $\mathbb{F}_{q}[x;\theta,\Im]$. In order to make Table 1 precise, we enlist the coefficients of polynomials in decreasing powers of $x$. For example, we write $w^{7}02w$ to represent the polynomial $w^{7}x^{3}+2x+w$.
5.3 Theory Behind the Encoding and Error Correction Procedure
5.3.1 Encoding
The dimension of the quantum code is $q^{2k-(s+1)n}$; hence, the remaining dimension $q^{2(s+1)n-2k}$ corresponds to the redundancy. Let |$\ket{\phi^{\prime}}$ be the $(2k-(s+1)n)$ qudit message state. We consider $(2(s+1)n-2k)$ qudits in state |$\ket{0}$ each called the ancillary qudits or ancilla qudits that correspond to the redundancy that is added to the code. The encoding of the stabilizer quantum codes involves applying an operator $\mathcal{E}$ to the state |$\ket{\phi^{\prime}}\ket{0}^{\otimes(2(s+1)n-2k)}$. The encoding operator $\mathcal{E}$ is a product of operators from a group called the Clifford group. While working with basis operators, we need unitary operators that transform a basis operator to another basis operator, called the Clifford operators [15]. The Clifford operators transform every Pauli basis operator into a Pauli basis operator. The set of all Clifford operators forms the Clifford group that is generated by the discrete Fourier transform $(\text{DFT}_{q})$ operator, phase shift operator, and the addition $(\text{ADD}_{q})$ operator [15, 20].
5.3.2 Error Correction
Syndrome computation involves computing the syndrome based on the erroneous state $E\ket{\psi}$, where $E$ is an error that belongs to the Pauli basis $\mathcal{P}^{\otimes(s+1)n}$. We apply the syndrome computation operator that operates on $E\ket{\psi}$ along with $(2(s+1)n-2k)$ syndrome qudits in state |$\ket{0}$ to transform it to $E\ket{\psi}\ket{s}$. Using the syndrome state |$\ket{s}$ as the control and the codeword qudits as the target, the inverse error operation $E^{\dagger}$ is applied to obtain the codeword |$\ket{\psi}$.
We next discuss the syndrome computation and error correction procedure when the error $E$ does not belong to the Pauli basis. The error $E$ belongs to $\mathbb{C}^{q^{(s+1)n}\times q^{(s+1)n}}$ as it is an $(s+1)n$ qudit operator; hence, $E$ can be expressed in terms of the Pauli basis $\mathcal{P}^{\otimes(s+1)n}$ as the Pauli basis is a basis for $\mathbb{C}^{q^{(s+1)n}\times q^{(s+1)n}}$. Let
$E=\sum_{B\in\mathcal{P}^{\otimes(s+1)n}}a_{B}B,\text{ where }a_{B}\in\mathbb{C},$
this $\text{this implies}\leavevmode\nobreak\ \leavevmode\nobreak\ E\ket{\psi}=\left(\sum_{B\in\mathcal{P}^{\otimes(s+1)n}}a_{B}B\right)\ket{\psi}=\sum_{B\in\mathcal{P}^{\otimes(s+1)n}}a_{B}B\ket{\psi}.$ (17)
Let us introduce $(2(s+1)n-2k)$ syndrome qudits in state |$\ket{0}$, then, we obtain
$E\ket{\psi}\ket{0}^{\otimes(2(s+1)n-2k)}=\sum_{B\in\mathcal{P}^{\otimes(s+1)n}}a_{B}B\ket{\psi}\ket{0}^{\otimes(2(s+1)n-2k)}.$ (18)
For the basis error $B$, the syndrome |$\ket{s_{B}}$ is obtained based on the eigenvalues of the stabilizers with respect to $B\ket{\psi}$. Let $\mathscr{S}$ be the syndrome computation operator that transforms $B\ket{\psi}\ket{0}^{\otimes(2(s+1)n-2k)}$ to $B\ket{\psi}\ket{s_{B}}$. Then we operate $\mathscr{S}$ on $E\ket{\psi}\ket{0}^{\otimes(2(s+1)n-2k)}$, and obtain
$\mathscr{S}E\ket{\psi}\ket{0}^{\otimes(2(s+1)n-2k)}$ $=\mathscr{S}\left(\sum_{B\in\mathcal{P}^{\otimes(s+1)n}}a_{B}B\ket{\psi}\ket{0}^{\otimes(2(s+1)n-2k)}\right)$
$=\sum_{B\in\mathcal{P}^{\otimes(s+1)n}}a_{B}\mathscr{S}\left(B\ket{\psi}\ket{0}^{\otimes(2(s+1)n-2k)}\right),$
$=\sum_{B\in\mathcal{P}^{\otimes(s+1)n}}a_{B}B\ket{\psi}\ket{s_{B}}.$ (19)
As |$\ket{s_{B}}$s are of the form |$\ket{s_{1}}\ket{s_{2}}\dots\ket{s_{(2(s+1)n-2k)}}$, they are orthogonal states for correctable errors. Thus, on measuring these syndrome qudits, the measurement outcome is $s_{B}=[s_{1}\leavevmode\nobreak\ s_{2}\leavevmode\nobreak\ \dots\\ s_{(2(s+1)n-2k)}]$ for some $B$ with the post-measurement state being $B\ket{\psi}\ket{s_{B}}$. Also, using the syndrome $s_{B}$, the error is deduced, and the inverse error $B^{\dagger}$ is applied. Here, the syndrome qudits are discarded.
Alternatively, using control-based operations with the $(2(s+1)n-2k)$ syndrome qudits are control qudits and the codeword qudits as target qudits, the inverse error operator $B^{\dagger}$ is applied when the syndrome state is |$\ket{s_{B}}$. Thus, the errors that are not Pauli basis errors are also corrected. We conclude that if we can correct a subset of Pauli basis errors, then we can correct errors that can be expressed as a linear combination of these errors.
@@ 6 Conclusion
In this paper, we have constructed many quantum codes over a class of finite commutative non-chain rings $\mathscr{R}_{q,s}$, with better parameters than the codes available in recent literature. Particularly, we have obtained $(\gamma,\Delta)$-cyclic codes using a set of idempotents over $\mathscr{R}_{q,s}$ and established results on their algebraic structure. Towards the construction of quantum codes, a necessary and sufficient condition to contain their dual codes has been established. Finally, we have obtained many better quantum codes. We have concluded our work by discussing the encoding and error correction capacity of our proposed quantum codes. However, exploring applications in the quantum computations of these codes is still open as future research work.
@@ Acknowledgements
The first and second authors are thankful to the Department of Science and Technology (DST), Govt. of India, for financial support under CRG/2020/005927, vide Diary No. SERB/F/6780/ 2020-2021 dated 31 December 2020 and Ref No. DST/INSPIRE/03/2016/001445, respectively.
@@
Data Availability Statement: The authors declare that [the/all other] data supporting the findings of this study are available within the article. Any clarification may be requested from the corresponding author provided it is essential.